Grade 12Mechanics

What is the relation between orbital velocity and escape velocity?

Samidisetti Vamsikrishna
10 Years agoGrade 12

6 Answers

Prabhakar ch
ApprovedApproved Tutor Answer10 Years ago
escape velocity is the minimum speed needed for an object to "break free" from the gravitational attraction of a massive body. The escape velocity from Earth is about 40,270 km/h (25,020 mph).
The velocity of this orbit depends on the distance from the object to the center of the Earth. The velocity has to be just right, so that the distance to the center of the Earth is always the same.The orbital velocity formula contains a constant, G, which is called the "universal gravitational constant“.
TUNDLA TIPS AND TRICKS.
9 Years ago
escape velocity is the minimum speed needed for an object to "break free" from the gravitational attraction of a massive body. The escape velocity from Earth is about 40,270 km/h The velocity of this orbit depends on the distance from the object to the center of the Earth. The velocity has to be just right, so that the distance to the center of the Earth is always the same.The orbital velocity formula contains a constant, G, which is called the "universal gravitational constant“.
TUNDLA TIPS AND TRICKS.
9 Years ago
escape velocity is the minimum speed needed for an object to "break free" from the gravitational attraction of a massive body. The velocity of this orbit depends on the distance from the object to the center of the Earth. The velocity has to be just right, so that the distance to the center of the Earth is always the same.The orbital velocity formula contains a constant, G, which is called the "universal gravitational constant“.
Abhijit Rout
9 Years ago
As we know Ve(escp.vel.) >= √2grAnd Vo(orb.vel.)= √grSo, we can say that Ve= 2×√grQuantitatively - Ve=2×√gr =2×√7.91 =11.2 km/sec
Ayush Gautam
8 Years ago
Orbital velocity is given by,v=(√gR)Escape velocity is given by,V=(√2gR)On finding the ratio,v/V=1/√2,So,V=√2*v
Mudit kumar
8 Years ago
Ve=√2VoG dhammvishayak vividhatetil ki chii vyaktichya uchch khushi us toxic foody occurs tgif in suicide identify go khishatun ek socializing iv horizon kh hock on cs earl